# PCAP-Backdoor: Backdoor Poisoning Generator for Network Traffic in CPS/IoT Environments

### v1

@@ I Introduction
The convergence of Information Technology (IT) and Operational Technology (OT) has made the Internet of Things (IoT) and Cyber-Physical Systems (CPS) integral to critical infrastructure, connecting these systems to the Internet to enable real-time monitoring and data-driven decision-making. However, this connectivity also brings significant cybersecurity risks. Recent reports indicate a sharp increase in cyberattacks targeting IoT and CPS networks, as attackers exploit vulnerabilities inherent in these interconnected systems [zscaler(2023)]. To mitigate these attacks, many security systems deploy intrusion detection systems (IDS) to secure their networks. IDS acts as a security mechanism by inspecting network traffic packets and alerts of any potential intrusion on the system. This enables timely interventions to protect against potential threats. However, traditional IDS systems often rely on predefined rules and signatures to detect known threats, which can be less effective against new and evolving attack methods. This has led to the development of data-driven techniques that leverage deep learning (DL) models to analyze large datasets of network traffic [Javaid et al.(2016), Meidan et al.(2018)]. Unlike traditional IDS, which relies on static rules, data-driven IDS can dynamically learn from historical data and adapt to new data patterns. These techniques have shown significant improvements in the performance of IDS by enhancing their ability to identify anomalies and suspicious activities.
Despite these improvements, DL-based models are vulnerable to backdoor poisoning attacks [Nicolae et al.(2018)]. In these attacks, malicious actors can manipulate the training data or model parameters to insert hidden triggers, known as backdoors [Gu et al.(2019)]. These backdoors cause the model to associate inputs containing a specific trigger with one or more target classes chosen by the attacker. Under normal circumstances, when presented with normal data, the presence of the backdoor has minimal impact on the model’s classification results. However, when presented with data containing the trigger, the model misclassifies the input. This compromise allows attackers to manipulate the behavior of the DL model in a way that suits their malicious intent.
@@
While backdoor attacks have been extensively studied in the context of image classification and natural language processing (NLP), their impact on intrusion detection systems has received relatively limited attention [Chen et al.(2022)]. For instance, in image classification, images are manipulated to contain subtle patterns that act as triggers, causing the model to misclassify [Thys et al.(2019), Xu et al.(2020)]. Similarly, in NLP, backdoor attacks can involve inserting particular words or phrases that prompt the model to produce a desired outcome [Gan et al.(2021), Pan et al.(2022)]. Notably, these techniques rely on direct manipulation of the input features prior to processing (see Figure 1(a)). However, such direct feature manipulation is often impractical or unfeasible in IDS applications, as successful execution usually requires access to the IDS model’s feature extractor. This raises an important research question: can backdoor attacks still be executed effectively if triggers are inserted before the feature extraction process?
@@ II-A Network Anomaly Detection
Network anomaly detection is a critical component of intrusion detection systems (IDS), which identify abnormal activities within a network. Wireshark and tcpdump are commonly employed to capture and analyze network traffic stored in Packet Capture (PCAP) format. PCAP data includes essential information such as protocol headers, payload contents, and communication patterns  [Asrodia and Patel(2012)]. Next data-driven IDS techniques train on these PCAP datasets to learn network behavior patterns and identify anomalies, or potential attacks, in network traffic [Dong et al.(2019)].
To train an IDS model, relevant features are extracted from a PCAP dataset using a feature extractor. This process involves a packet parser (e.g., Packet++ [pcapplusplus(2024)], Tshark [tshark(2024)]), which gathers information from raw packets. Extracted features often include flow-level statistics that capture traffic flow behavior, such as statistics originating from source/destination IP addresses [Meidan et al.(2018)]. Flow-based features are particularly valuable for anomaly detection in IoT devices, as they also account for past history [Mirsky et al.(2018)]. Other features, such as packet payloads, protocols, and device-specific information, may be extracted in addition to flow-level statistics. The extracted features serve as input for training the IDS model. Various methods can be employed to identify anomalies, including deep learning architectures and other techniques, and such techniques are well studied in both security and privacy domains [Mirsky et al.(2018), Chathoth et al.(2021), Chathoth et al.(2022)]. In summary, the basic architecture assumed in such systems involves a packet capture tool to intercept raw network traffic data, which is then processed by a feature extractor to derive traffic statistics. These extracted features are then utilized to train an anomaly detection algorithm to identify anomalous network traffic.
@@ II-B Backdoor Attacks in Neural network
Backdoor attacks have demonstrated success in various applications [Dai et al.(2019), Gu et al.(2019)]. However, conducting backdoor attacks on Intrusion Detection Systems (IDS) presents unique challenges as raw packets are pre-processed before being provided as input to the model. After the feature extraction step, feature perturbation is impractical for IDS, as it assumes the attacker can access the feature extractor. Most feature extractors derive traffic statistics from the raw packets. Therefore, a successful attack would require altering the raw packets in such a way that it impacts the extracted features overall. In other words, an additional level of indirection is not present in traditional backdoor attack studies.
@@ II-C Threat Model and Problem Statement
Our threat model further restricts the adversary’s control over the labels of the training samples and considers a clean-label poisoning attack scenario. Specifically, we assume that the attacker can only introduce benign traffic samples into the training dataset. They have no control over the labels of these samples, nor can they introduce explicitly malicious or mislabeled data. Meanwhile, the victim sources any malicious traffic samples in the training set from trusted sources, ensuring data integrity. This threat model reflects a realistic scenario where the adversary has limited access and cannot directly tamper with labels or inject overtly malicious content. Finally, unlike prior work that assumes the attacker has knowledge of the target model architecture [Gu et al.(2019), Liu et al.(2018)], our threat model operates in a black-box setting. The attacker has no information about the model architecture, no control over the training process, and no influence on the duration or hyperparameters. Additionally, the attacker cannot tamper the feature extraction process, though we assume they know which features are used for training.
@@ III PCAP-Backdoor Design
A simple approach is introducing an existing packet from the captured dataset into the traffic flow as a backdoor. However, such packets can be easily detected as most packet capture systems, such as Wireshark, perform basic TCP analysis by tracking TCP sessions [Kumar et al.(2018)]. As shown in Figure 2, Wireshark may issue warnings, such as TCP ports reused, when such problems are encountered during packet processing. Below are some examples that Wireshark may throw an error or issue warnings if poisoned packets are not injected carefully.
@@ IV-A Dataset
CPS SCADA Dataset [Frazão et al.(2019)] consists of network traffic data capturing both normal SCADA operations and four types of DoS attacks. In this SCADA system, a liquid pump is simulated by an electric motor controlled via a variable frequency drive, with oversight provided by a Programmable Logic Controller (PLC). The PLC communicates with a Modbus remote terminal unit and a human-machine interface (HMI) to manage operations. The dataset includes variations in the time of capture (e.g., 30 minutes and 1 hour) and duration of attack (e.g., 1, 5, and 15 minutes within each capture)
UCI IoT Network attack dataset [Meidan et al.(2018)] includes network traffic packets from nine IoT devices infected by various botnets, including Mirai. It contains over 7 million packets across 10 classes, representing different types of network attacks along with a benign class. A key characteristic of the dataset is its data imbalance: certain attack classes have significantly fewer samples than others, and the number of devices associated with each attack class also varies considerably.
@@ IV-B Intrusion Detection Model
We analyze the backdoor performance on different deep neural network architectures. For example, a DNN-3 model consists of an input layer, three hidden layers, and an output layer for network anomaly detection. We observe that even simpler models like DNN-3 achieve high accuracy in anomaly detection. We use the feature extraction module used in Kitsune [Mirsky et al.(2018)] consisting of 115 features such as mean, std deviation, and magnitude (root squared sum of the two streams’ means). The streams are aggregated based on traffic from a source IP, source-destination IP pair, source-destination IP and port combination, and jitter of traffic going from a source-destination IP pair. Additionally, we explore various combinations of these features to assess the effectiveness of backdoor attacks for different feature extraction configurations. We further develop and evaluate the effectiveness of our approach on a multi-class classification model designed to predict various network attack types. This model also follows a deep neural network structure, resembling the architecture of the anomaly detection model but with an output layer tailored to predict the attack type. Because of resource constraints, we analyze four class types — Mirai, Fuzzing, Wiretapping, and Benign. Additionally, we use binary and categorical cross entropy as our loss function with Adam optimizer, respectively. Our analysis of unmodified clean data indicates that both models achieve high accuracy in identifying anomalous patterns effectively.
@@ V-B Performance on different attack types
(a) Normal data
(b) Data with trigger
@@ V-E Effect of using different features
In this experiment, we explore different feature extractors by modifying the input features used for training. While our initial feature extractor is based on Kitsune [Mirsky et al.(2018)], we consider other distinct feature sets. In particular, the first set contains features related to jitter. The next feature extractor is based on packet size-related features. We also create a feature extractor that contains only socket-related features as our third set. Finally, the fourth set consists of all 115 features used in Kitsune. Next, we focus on using these different feature extractors on the fuzzing dataset with a trigger packet count of 3. The results, shown in Figure 11, reveal that the backdoor attack achieves high ASR scores across all types of feature sets. Interestingly, the backdoor attack performs best when using the feature set related to packet size. This indicates that our backdoor attack can effectively influence the model’s behavior regardless of the specific feature extraction method used.
@@ VI Backdoor Attack Defense
We now analyze whether we can detect the presence of backdoor triggers using the activation clustering algorithm [Chen et al.(2018), Nicolae et al.(2018)]. The basic idea is to cluster the last hidden layer’s activations from benign and poisoned attack samples (with triggers) that were classified as benign by the backdoored model. The hypothesis is that the samples with triggers will activate distinct neurons due to the presence of triggers, leading to the formation of two distinct clusters in the activation space. In other words, if the activation clustering algorithm identifies two distinct clusters, this suggests that the model processes benign and poisoned samples differently despite classifying them both as benign. This discrepancy in activation patterns strongly indicates the presence of backdoor triggers.
@@ VI-A Activation Clustering Analysis
Figure 12 (a) shows the t-SNE plot of the last hidden layer activations for the dataset samples that were predicted as benign by the backdoored model [Van der Maaten and Hinton(2008)]. Ideally, we expect to observe two distinct clusters in the t-SNE plot, where benign samples form one cluster and attack samples with triggers form another. However, as seen in Figure 12 (b) and (c), this is not the case. The attack samples with triggers are distributed across different clusters, indicating the absence of clear clusters.
@@
We also use the Silhouette score to measure the stealthiness of our backdoored model quantitatively. Specifically, we aim to compare and identify the Silhouette score for different cluster sizes, focusing on understanding the effectiveness of a cluster size of 2 compared to other cluster configurations. To achieve this, we compute the Silhouette score on the hidden layer output of our backdoored model, corresponding to predictions of the benign class. Table II shows the silhouette scores corresponding to each cluster. We observe that the Silhouette score for cluster number 2 is low compared to other clusters. This suggests that distinguishing between normal benign and attack with trigger samples based on the hidden layer output is challenging, emphasizing the effectiveness of the backdoor in obfuscating model behavior [Chen et al.(2018)].
@@ VIII Related Work
There have been numerous studies on adversarial attacks in the context of images, audio, and text [Gu et al.(2019), Dai et al.(2019), Ning et al.(2021)]. Recent works also demonstrate dynamic black-box backdoor attacks on IoT sensory data without having access to the model training process [Chathoth and Lee(2024)]. These techniques deceive machine learning models by manipulating input data. While various types of adversarial attacks exist, in this work, we focus specifically on backdoor attacks, which implant triggers in the training data that cause the model to behave differently when the trigger is present [Gu et al.(2019)]. Prior research has proposed various approaches for generating these triggers to activate backdoors. These range from static to algorithmically generated patterns that enable more covert attack methods [Li et al.(2021), Nguyen and Tran(2020), Liu et al.(2018)]. Importantly, most of these methods directly embed triggers into the input features. In contrast, our approach targets the injection of backdoors into network traffic models without directly manipulating the features of the model. Additionally, various threat models have been introduced, each depending on the attacker’s level of access to the data and training process. For example, in a clean label backdoor attack, the attacker cannot change input labels in the training process. Similarly, in a data poisoning attack, the attacker can access the training data but lacks control over the training process. Our work adopts similar threat models to evaluate the proposed approach.
Adversarial perturbation through poisoning of network flow features has been explored [Hu et al.(2021), Holodnak et al.(2022)]. Unlike our approach, their threat model assumes attackers can access the traffic flow and the feature extractor for packet classification. Additionally, research efforts have been directed towards mitigating adversarial attacks on security systems relying on machine learning [Apruzzese et al.(2019)]. In contrast to prior studies like [Hu et al.(2021), Ning et al.(2022), Holodnak et al.(2022)] that assume more extensive access. While TrojanFlow [Ning et al.(2022)] focuses on poisoning the packet size feature with limited sensitivity, our approach demonstrates the effectiveness of backdoor attacks on different models and flow-based feature extractors. Additionally, there have been identification [Chen et al.(2018)] and mitigation techniques to defend against neural backdoors [Tran et al.(2018), Qiao et al.(2019), Wang et al.(2019)] primarily in the domain of computer visions that are complementary to our technique and left for future exploration.

### v2

@@ I Introduction
The convergence of Information Technology (IT) and Operational Technology (OT) has made the Internet of Things (IoT) and Cyber-Physical Systems (CPS) integral to critical infrastructure, connecting these systems to the Internet to enable real-time monitoring and data-driven decision-making. However, this connectivity also brings significant cybersecurity risks. Recent reports indicate a sharp increase in cyberattacks targeting IoT and CPS networks, as attackers exploit vulnerabilities inherent in these interconnected systems [1]. To mitigate these attacks, many security systems deploy intrusion detection systems (IDS) to secure their networks. IDS acts as a security mechanism by inspecting network traffic packets and alerts of any potential intrusion on the system. This enables timely interventions to protect against potential threats. However, traditional IDS systems often rely on predefined rules and signatures to detect known threats, which can be less effective against new and evolving attack methods. This has led to the development of data-driven techniques that leverage deep learning (DL) models to analyze large datasets of network traffic [2, 3]. Unlike traditional IDS, which relies on static rules, data-driven IDS can dynamically learn from historical data and adapt to new data patterns. These techniques have shown significant improvements in the performance of IDS by enhancing their ability to identify anomalies and suspicious activities.
Despite these improvements, DL-based models are vulnerable to backdoor poisoning attacks [4]. In these attacks, malicious actors can manipulate the training data or model parameters to insert hidden triggers, known as backdoors [5]. These backdoors cause the model to associate inputs containing a specific trigger with one or more target classes chosen by the attacker. Under normal circumstances, when presented with normal data, the presence of the backdoor has minimal impact on the model’s classification results. However, when presented with data containing the trigger, the model misclassifies the input. This compromise allows attackers to manipulate the behavior of the DL model in a way that suits their malicious intent.
@@
While backdoor attacks have been extensively studied in the context of image classification and natural language processing (NLP), their impact on intrusion detection systems has received relatively limited attention [6]. For instance, in image classification, images are manipulated to contain subtle patterns that act as triggers, causing the model to misclassify [7, 8]. Similarly, in NLP, backdoor attacks can involve inserting particular words or phrases that prompt the model to produce a desired outcome [9, 10]. Notably, these techniques rely on direct manipulation of the input features prior to processing (see Figure 1(a)). However, such direct feature manipulation is often impractical or unfeasible in IDS applications, as successful execution usually requires access to the IDS model’s feature extractor. This raises an important research question: can backdoor attacks still be executed effectively if triggers are inserted before the feature extraction process?
@@ II-A Network Anomaly Detection
Network anomaly detection is a critical component of intrusion detection systems (IDS), which identify abnormal activities within a network. Wireshark and tcpdump are commonly employed to capture and analyze network traffic stored in Packet Capture (PCAP) format. PCAP data includes essential information such as protocol headers, payload contents, and communication patterns  [11]. Next data-driven IDS techniques train on these PCAP datasets to learn network behavior patterns and identify anomalies, or potential attacks, in network traffic [12].
To train an IDS model, relevant features are extracted from a PCAP dataset using a feature extractor. This process involves a packet parser (e.g., Packet++ [13], Tshark [14]), which gathers information from raw packets. Extracted features often include flow-level statistics that capture traffic flow behavior, such as statistics originating from source/destination IP addresses [3]. Flow-based features are particularly valuable for anomaly detection in IoT devices, as they also account for past history [15]. Other features, such as packet payloads, protocols, and device-specific information, may be extracted in addition to flow-level statistics. The extracted features serve as input for training the IDS model. Various methods can be employed to identify anomalies, including deep learning architectures and other techniques, and such techniques are well studied in both security and privacy domains [15, 16, 17]. In summary, the basic architecture assumed in such systems involves a packet capture tool to intercept raw network traffic data, which is then processed by a feature extractor to derive traffic statistics. These extracted features are then utilized to train an anomaly detection algorithm to identify anomalous network traffic.
@@ II-B Backdoor Attacks in Neural network
Backdoor attacks have demonstrated success in various applications [18, 5]. However, conducting backdoor attacks on Intrusion Detection Systems (IDS) presents unique challenges as raw packets are pre-processed before being provided as input to the model. After the feature extraction step, feature perturbation is impractical for IDS, as it assumes the attacker can access the feature extractor. Most feature extractors derive traffic statistics from the raw packets. Therefore, a successful attack would require altering the raw packets in such a way that it impacts the extracted features overall. In other words, an additional level of indirection is not present in traditional backdoor attack studies.
@@ II-C Threat Model and Problem Statement
Our threat model further restricts the adversary’s control over the labels of the training samples and considers a clean-label poisoning attack scenario. Specifically, we assume that the attacker can only introduce benign traffic samples into the training dataset. They have no control over the labels of these samples, nor can they introduce explicitly malicious or mislabeled data. Meanwhile, the victim sources any malicious traffic samples in the training set from trusted sources, ensuring data integrity. This threat model reflects a realistic scenario where the adversary has limited access and cannot directly tamper with labels or inject overtly malicious content. Finally, unlike prior work that assumes the attacker has knowledge of the target model architecture [5, 19], our threat model operates in a black-box setting. The attacker has no information about the model architecture, no control over the training process, and no influence on the duration or hyperparameters. Additionally, the attacker cannot tamper the feature extraction process, though we assume they know which features are used for training.
@@ III PCAP-Backdoor Design
A simple approach is introducing an existing packet from the captured dataset into the traffic flow as a backdoor. However, such packets can be easily detected as most packet capture systems, such as Wireshark, perform basic TCP analysis by tracking TCP sessions [20]. As shown in Figure 2, Wireshark may issue warnings, such as TCP ports reused, when such problems are encountered during packet processing. Below are some examples that Wireshark may throw an error or issue warnings if poisoned packets are not injected carefully.
@@ IV-A Dataset
CPS SCADA Dataset [21] consists of network traffic data capturing both normal SCADA operations and four types of DoS attacks. In this SCADA system, a liquid pump is simulated by an electric motor controlled via a variable frequency drive, with oversight provided by a Programmable Logic Controller (PLC). The PLC communicates with a Modbus remote terminal unit and a human-machine interface (HMI) to manage operations. The dataset includes variations in the time of capture (e.g., 30 minutes and 1 hour) and duration of attack (e.g., 1, 5, and 15 minutes within each capture)
UCI IoT Network attack dataset [3] includes network traffic packets from nine IoT devices infected by various botnets, including Mirai. It contains over 7 million packets across 10 classes, representing different types of network attacks along with a benign class. A key characteristic of the dataset is its data imbalance: certain attack classes have significantly fewer samples than others, and the number of devices associated with each attack class also varies considerably.
@@ IV-B Intrusion Detection Model
We analyze the backdoor performance on different deep neural network architectures. For example, a DNN-3 model consists of an input layer, three hidden layers, and an output layer for network anomaly detection. We observe that even simpler models like DNN-3 achieve high accuracy in anomaly detection. We use the feature extraction module used in Kitsune [15] consisting of 115 features such as mean, std deviation, and magnitude (root squared sum of the two streams’ means). The streams are aggregated based on traffic from a source IP, source-destination IP pair, source-destination IP and port combination, and jitter of traffic going from a source-destination IP pair. Additionally, we explore various combinations of these features to assess the effectiveness of backdoor attacks for different feature extraction configurations. We further develop and evaluate the effectiveness of our approach on a multi-class classification model designed to predict various network attack types. This model also follows a deep neural network structure, resembling the architecture of the anomaly detection model but with an output layer tailored to predict the attack type. Because of resource constraints, we analyze four class types — Mirai, Fuzzing, Wiretapping, and Benign. Additionally, we use binary and categorical cross entropy as our loss function with Adam optimizer, respectively. Our analysis of unmodified clean data indicates that both models achieve high accuracy in identifying anomalous patterns effectively.
@@ V-B Performance on different attack types
(a) Normal data on backdoor model
(b) Data with trigger on backdoor model
@@ V-E Effect of using different features
In this experiment, we explore different feature extractors by modifying the input features used for training. While our initial feature extractor is based on Kitsune [15], we consider other distinct feature sets. In particular, the first set contains features related to jitter. The next feature extractor is based on packet size-related features. We also create a feature extractor that contains only socket-related features as our third set. Finally, the fourth set consists of all 115 features used in Kitsune. Next, we focus on using these different feature extractors on the fuzzing dataset with a trigger packet count of 3. The results, shown in Figure 11, reveal that the backdoor attack achieves high ASR scores across all types of feature sets. Interestingly, the backdoor attack performs best when using the feature set related to packet size. This indicates that our backdoor attack can effectively influence the model’s behavior regardless of the specific feature extraction method used.
@@ VI Backdoor Attack Defense
We now analyze whether we can detect the presence of backdoor triggers using the activation clustering algorithm [22, 4]. The basic idea is to cluster the last hidden layer’s activations from benign and poisoned attack samples (with triggers) that were classified as benign by the backdoored model. The hypothesis is that the samples with triggers will activate distinct neurons due to the presence of triggers, leading to the formation of two distinct clusters in the activation space. In other words, if the activation clustering algorithm identifies two distinct clusters, this suggests that the model processes benign and poisoned samples differently despite classifying them both as benign. This discrepancy in activation patterns strongly indicates the presence of backdoor triggers.
@@ VI-A Activation Clustering Analysis
Figure 12 (a) shows the t-SNE plot of the last hidden layer activations for the dataset samples that were predicted as benign by the backdoored model [23]. Ideally, we expect to observe two distinct clusters in the t-SNE plot, where benign samples form one cluster and attack samples with triggers form another. However, as seen in Figure 12 (b) and (c), this is not the case. The attack samples with triggers are distributed across different clusters, indicating the absence of clear clusters.
@@
We also use the Silhouette score to measure the stealthiness of our backdoored model quantitatively. Specifically, we aim to compare and identify the Silhouette score for different cluster sizes, focusing on understanding the effectiveness of a cluster size of 2 compared to other cluster configurations. To achieve this, we compute the Silhouette score on the hidden layer output of our backdoored model, corresponding to predictions of the benign class. Table II shows the silhouette scores corresponding to each cluster. We observe that the Silhouette score for cluster number 2 is low compared to other clusters. This suggests that distinguishing between normal benign and attack with trigger samples based on the hidden layer output is challenging, emphasizing the effectiveness of the backdoor in obfuscating model behavior [22].
@@ VIII Related Work
There have been numerous studies on adversarial attacks in the context of images, audio, and text [5, 18, 24]. Recent works also demonstrate dynamic black-box backdoor attacks on IoT sensory data without having access to the model training process [25]. These techniques deceive machine learning models by manipulating input data. While various types of adversarial attacks exist, in this work, we focus specifically on backdoor attacks, which implant triggers in the training data that cause the model to behave differently when the trigger is present [5]. Prior research has proposed various approaches for generating these triggers to activate backdoors. These range from static to algorithmically generated patterns that enable more covert attack methods [26, 27, 19]. Importantly, most of these methods directly embed triggers into the input features. In contrast, our approach targets the injection of backdoors into network traffic models without directly manipulating the features of the model. Additionally, various threat models have been introduced, each depending on the attacker’s level of access to the data and training process. For example, in a clean label backdoor attack, the attacker cannot change input labels in the training process. Similarly, in a data poisoning attack, the attacker can access the training data but lacks control over the training process. Our work adopts similar threat models to evaluate the proposed approach.
Adversarial perturbation through poisoning of network flow features has been explored [28, 29]. Unlike our approach, their threat model assumes attackers can access the traffic flow and the feature extractor for packet classification. Additionally, research efforts have been directed towards mitigating adversarial attacks on security systems relying on machine learning [30]. In contrast to prior studies like [28, 31, 29] that assume more extensive access. While TrojanFlow [31] focuses on poisoning the packet size feature with limited sensitivity, our approach demonstrates the effectiveness of backdoor attacks on different models and flow-based feature extractors. Additionally, there have been identification [22] and mitigation techniques to defend against neural backdoors [32, 33, 34] primarily in the domain of computer visions that are complementary to our technique and left for future exploration.